help me please!
find the measure of the area of the shaded region. Simplify.
inner rectangle:
L= 2x
W= 8x^2 + 12x + 8
outer rectangle:
L= 5x
W= 12x^2 + 9x - 6

Answer by vasumathi(46)

Solution:inner rectangle:
L= 2x
W= 8x^2 + 12x + 8
Area of the inner rectangle is ..
L*W = 2x(8x^2+12x +8)=16 x^3+24 x^2+16x
outer rectangle:
L= 5x
W= 12x^2 + 9x - 6
Area of the outer rectangle is ...
L*W = 5x(12x^2+9x-6)=60 x^3+45x^2-30x
The area of the shaded portion is the difference of these two areas
A = 60 x^3+45x^2-30x-(16 x^3+24 x^2+16x )
A = 60 x^3+45x^2-30x-16x^3-24x^2-16x
A = 44 x^3 +21 x^2-46 x